# SOOD++: Leveraging Unlabeled Data to Boost Oriented Object Detection

### v1

@@ 5.1 Dataset and Evaluation Metric
TABLE I: The detailed hyper-parameters of our method.
     Hyper-parameter      Value
     $\psi$      50
     Sample ratio $\mathcal{R}$      0.25
     IoU threshold $\mathcal{T}_{h}$      0.1
     Noisy intensity $\beta$      0.3
     Global number threshold $\mathcal{T}_{g}$      150
DOTA-V1.5 [58] contains 2,806 large aerial images and 402,089 annotated oriented objects. It includes three subsets: DOTA-V1.5-train, DOTA-V1.5-val, and DOTA-V1.5-test, containing 1,411, 458, and 937 images, respectively. The annotations of the DOTA-V1.5-test are not released. There are 16 categories in this dataset. Compared with its previous version, DOTA-V1.0, DOTA-V1.5 contains more small instances (less than 10 pixels), which makes it more challenging. For the partially labeled data setting, we randomly sample 10%, 20%, and 30% of the images from DOTA-V1.5-train as labeled data and set the remaining images as unlabeled data. For the fully labeled data setting, we set DOTA-V1.5-train as labeled data and DOTA-V1.5-test as unlabeled data. We report the performance with the standard mean average precision (mAP).
@@
DOTA-V1.0 [58] contains the same images and data splits as DOTA-V1.5, including 1,411 training images, 458 validation images, and 937 test images. As an early version of the DOTA-series datasets, it contains 15 categories identical to DOTA-V1.5, except for annotating container-crane. For the partially labeled data setting, we use the same labeled and unlabeled splits as in DOTA-V1.5, including 10%, 20%, and 30%. For the fully labeled data setting, we set DOTA-V1.0-train as labeled data and DOTA-V1.0-test as unlabeled. The mAP is used as the evaluation metric.
@@
HRSC2016 [41] is a widely-used dataset for arbitrarily oriented ship detection, consisting of 436 training images, 181 validation images, and 444 test images, with sizes ranging from 300×300 to 1,500×900. For the partially labeled data setting, we randomly sample 10%, 20%, and 30% images from the training set as labeled data and set the remaining images as unlabeled data. For the fully labeled data setting, we set the training set as labeled data and the test set as unlabeled data. We report the AP85 and mAP50:95 performance on the validation set.
@@
The nuScenes Dataset [2] is a large-scale oriented 3D object detection dataset containing 1,000 distinct scenarios and 1.4M oriented 3D bounding boxes from 10 categories. These scenarios are recorded with a 32-beam LiDAR, six surround-view cameras, and five radars. The scenes are officially split into 700/150/150 scenes for the training/validation/test set. Each scene has 20s video frames and is fully annotated with 3D bounding boxes every 0.5 seconds. For the partially labeled data setting, we randomly sample 1/12, 1/6, and 1/3 annotated frames from the training set as labeled data and set the remaining training frames as unlabeled data. For the fully labeled data setting, we set all annotated frames from the training set as labeled data and unannotated frames as unlabeled data. We use the mAP and NDS as the primary metrics. Higher values indicate better performance, and the detailed information can be found in  [2].
@@ 5.2 Implementation Details
For the DOTA-V1.5, DOTA-V1.0, and HRSC2016 datasets, we take FCOS [56] as the representative anchor-free detector (the ResNet-50 [22] with FPN [33] used as the backbone), where BEVDet [24] is adopted for the nuScenes dataset. Following previous works [58, 20, 21], we crop the original images into 1,024$\times$1,024 patches with a stride of 824. This means the pixel overlap between two adjacent patches is 200 for the DOTA series datasets. We utilize asymmetric data augmentation for unlabeled data. We use strong augmentation (random flipping, color jittering, grayscale, and Gaussian blur) for the student model and weak augmentation (random flipping) for the teacher model. The models are trained for 180k iterations on 2 $\times$ RTX3090 GPUs (except the nuScenes dataset, using 8 $\times$ V100). With the SGD optimizer, the initial learning rate of 0.0025 is divided by ten at 120k and 160k. The momentum and the weight decay are set to 0.9 and 0.0001, respectively. Each GPU takes three images as input, where the proportion between unlabeled and labeled data is set to 1:2. Following previous SSOD works [76, 39], we use the “burn-in” strategy to initialize the teacher model. The detailed hyper-parameters are shown in Tab. I.
TABLE II: Experimental results on the DOTA-V1.5 (val set) with single-scale training and testing under 10%, 20%, and 30% partially labeled data. and indicate implementations with rotated-Faster R-CNN and rotated-FCOS, respectively. Colored numbers show performance improvement over the previous conference version (also one of the SOTA methods).
@@
Supervised-baseline Rotated-Faster R-CNN [50] NeurIPS 2016 43.43 51.32 53.14
Rotated-FCOS [56] ICCV 2019 42.78 50.11 54.79
Semi-supervised Unbiased Teacher [39] ICLR 2021 44.51 52.80 53.33
Soft Teacher [64] ICCV 2021 48.46 54.89 57.83
Dense Teacher [76] ECCV 2022 46.90 53.93 57.86
ARSL [35] CVPR 2023 47.56 55.05 58.94
SOOD [23] (conference version) CVPR 2023 48.63 55.58 59.23
SOOD++ (ours) - 50.48(+1.85) 57.44(+1.86) 61.51(+2.28)
TABLE III: Experimental results on DOTA-V1.5 (val set) with single-scale training and testing under fully labeled data. Numbers before the arrow are the supervised baseline ( for rotated-Faster R-CNN and for rotated-FCOS).
Method Publication mAP
Unbiased Teacher [39] ICLR 2021 66.12 $\xrightarrow{-1.27}$ 64.85
Soft Teacher [64] ICCV 2021 66.12 $\xrightarrow{+0.28}$ 66.40
Dense Teacher [76] ECCV 2022 65.46 $\xrightarrow{+0.92}$ 66.38
ARSL [35] CVPR 2023 65.46 $\xrightarrow{+1.22}$ 66.68
SOOD++ (ours) - 65.46 $\xrightarrow{+3.58}$ 69.04
6 Results and Analysis
In this section, we compare our method with the state-of-the-art SSOD methods [76, 64, 39, 35] on several challenging datasets. For a fair comparison, we re-implement these methods and have carefully tuned their hyper-parameters to obtain their best performance on oriented object detectors with the same augmentation setting.
6.1 Main Results
6.1.1 Results on the DOTA-V1.5 dataset
We first evaluate our method on the large-scale, challenging, and most convincing oriented object detection dataset, i.e., the DOTA-V1.5 dataset.
Partially labeled data. We conduct a comprehensive evaluation of our SOOD++ across various proportions of labeled data, as shown in Tab. II. We can make the following observations: 1) Our SOOD++ achieves state-of-the-art performance under all proportions, achieving 50.48, 57.44, and 61.51 mAP on 10%, 20%, and 30% proportions, respectively, significantly surpassing our supervised baseline rotated-FOCS [56] by 7.70, 7.33, and 6.72 mAP. 2) Compared with the state-of-the-art anchor-free method ARSL [35] and our conference version [23], we significantly outperform them by 2.92/1.85, 2.39/1.86 and 2.57/2.28 mAP, highly demonstrating the substantial improvements we made over the conference version. 3) Compared with Unbiased Teacher [39] and Soft Teacher [64], two representative anchor-based methods, on 10% and 20% proportions, our SOOD++ achieves higher performance even though our supervised baseline (rotated-FCOS) is weaker than they used (rotated-Faster R-CNN). Under 30% data proportion, SOOD++ surpasses Unbiased Teacher [39] and Soft Teacher [64] by a large margin, i.e., 8.18 and 3.68 mAP, respectively.
Fully labeled data. We further compare our SOOD++ with other semi-supervised object detection (SSOD) approaches under the fully labeled data setting, which is more challenging. The goal of this setting is to use additional unlabeled data to enhance the performance of a well-trained detector on large-scale labeled data. As shown in Tab. III, the results indicate that we surpass previous SSOD methods [76, 64, 39] by at least 2.36 mAP. Besides, compared with our baseline, even though it has been trained well in massive labeled data, we still achieve 3.58 mAP improvement, demonstrating our method’s ability to learn from unlabeled data. An interesting phenomenon is that the performance of Unbiased Teacher [39] drops after adding unlabeled data. We argue the reason might be that Unbiased Teacher does not apply unsupervised losses for bounding box regression, which is essential for oriented object detection.
Figure 6: Experimental results on the DOTA-V1.0 (val set) with single-scale training and testing under 10%, 20%, and 30% partially labeled data. All methods are implemented with rotated-FCOS.
@@ 6.1 Main Results
Supervised-baseline Rotated-FCOS [56] ICCV 2019 46.20 53.88 57.30
Semi-supervised Dense Teacher [76] ECCV 2022 50.87 58.02 60.91
ARSL [35] CVPR 2023 52.89 59.84 62.92
SOOD++ (ours) - 54.17 60.53 64.93
Figure 7: Experimental results on DOTA-V1.0 (val set) with single-scale training and testing under fully labeled data. Numbers before the arrow are the supervised baseline (rotated-FCOS).
Method Publication mAP
Dense Teacher [76] ECCV 2022 70.22 $\xrightarrow{+0.98}$ 71.20
ARSL [35] CVPR 2023 70.22 $\xrightarrow{+1.41}$ 71.63
SOOD++ (ours) - 70.22 $\xrightarrow{+2.18}$ 72.40
TABLE IV: Experimental results on HRSC2016 (val set) with single-scale training and testing under 10%, 20%, and 30% partially labeled data. All methods are implemented with rotated-FCOS.
@@
AP85(%) mAP50:90(%) AP85(%) mAP50:90(%) AP85(%) mAP50:90(%)
Supervised Rotated-FCOS [56] ICCV 2019 7.30 37.49 30.20 52.57 37.40 58.78
Semi-supervised Dense Teacher [76] ECCV 2022 21.90 49.93 32.80 59.63 41.40 63.08
ARSL [35] CVPR 2023 21.50 49.17 33.60 59.76 41.10 63.35
SOOD++ (ours) - 22.90 52.62 40.30 62.14 47.80 65.27
6.1.2 Results on the DOTA-V1.0 dataset
We then conduct the experiments on the DOTA-V1.0 dataset, which has the same images as the DOTA-V1.5 dataset but provides coarse annotations (e.g., many tiny objects with fewer pixels are not annotated).
Partially labeled data. We evaluate our method under different labeled data proportions on the DOTA-V1.0 dataset, as shown in Tab. 7. It is similar to the comparison of the DOTA-V1.5 dataset, where our SOOD++ achieves state-of-the-art performance under all proportions. Specifically, it surpasses the SOTA method [35] by 1.28, 0.69, and 2.01 mAP under 10%, 20%, and 30% labeled data settings, respectively. These impressive results demonstrate our method handles the coarse annotation dataset well.
Fully labeled data. Under this setting, we can explore the potential of the proposed semi-supervised framework even if it has been trained with extensive labeled data. As shown in Tab. 7, although the supervised baseline has already achieved high performance, our SOOD++ still takes 2.18 mAP improvements on it, an observed performance gain when using the unlabeled data. In addition, our method performs better compared with other semi-supervised methods [76, 35] with the same baseline.

### v2

@@ 5.1 Dataset and Evaluation Metric
TABLE I: Comparison on the DOTA-V2.0 (val set) under 10%, 20%, and 30% partially labeled data settings, using single-scale training and testing. // indicate implementations using rotated Faster R-CNN/FCOS/Deformable DETR.
Setting Method Publication 10% 20% 30%
Supervised-baseline Rotated-Faster R-CNN [32] NeurIPS 15 39.72 47.58 49.77
Rotated-FCOS [56] ICCV 19 38.78 46.74 49.62
Rotated Deformable DETR [69] ICLR 21 40.03 47.86 50.27
Semi-supervised Unbiased Teacher [1] ICLR 21 41.06 48.79 49.75
Soft Teacher [2] ICCV 21 44.97 50.68 53.23
Dense Teacher [3] ECCV 22 42.73 49.51 53.17
ARSL [5] CVPR 23 43.38 51.13 53.93
TMR-RD-v2 [7] WACV 24 45.65 51.80 54.24
Sparse Semi-DETR [6] CVPR 24 46.24 52.11 54.15
SOOD [53] CVPR 23 46.99 52.35 54.57
SOOD++(ours) - 49.14(+2.15) 54.27(+1.92) 56.90(+2.33)
TABLE II: Comparison on the DOTA-V2.0 (val set) with single-scale training/testing under fully labeled data. Results before the arrow are the supervised baseline (// for rotated Faster R-CNN/FCOS/Deformable DETR).
Method Publication 100%
Unbiased Teacher [1] ICLR 21 61.04 $\xrightarrow{-0.39}$ 60.65
Soft Teacher [2] ICCV 21 61.04 $\xrightarrow{+0.75}$ 61.79
Dense Teacher [3] ECCV 22 60.69 $\xrightarrow{+0.86}$ 61.55
ARSL [5] CVPR 23 60.69 $\xrightarrow{+1.37}$ 62.06
TMR-RD-v2 [7] WACV 24 61.04 $\xrightarrow{+0.71}$ 61.75
Sparse Semi-DETR [6] CVPR 24 60.73 $\xrightarrow{+0.78}$ 61.51
SOOD [53] CVPR 23 60.69 $\xrightarrow{+2.06}$ 62.75
SOOD++ (ours) - 60.69 $\xrightarrow{+3.62}$ 64.31
TABLE III: Comparison on the DOTA-V1.5 (val set) under 10%, 20%, and 30% partially labeled data settings, using single-scale training and testing. // indicate implementations using rotated Faster R-CNN/FCOS/Deformable DETR.
Setting Method Publication 10% 20% 30%
Supervised-baseline Rotated-Faster R-CNN [32] NeurIPS 15 43.43 51.32 53.14
Rotated-FCOS [56] ICCV 19 42.78 50.11 54.79
Rotated Deformable DETR [69] ICLR 21 43.55 51.41 53.85
Semi-supervised Unbiased Teacher [1] ICLR 21 44.51 52.80 53.33
Soft Teacher [2] ICCV 21 48.46 54.89 57.83
Dense Teacher [3] ECCV 22 46.90 53.93 57.86
ARSL [5] CVPR 23 47.56 55.05 58.94
TMR-RD-v2 [7] WACV 24 47.88 55.17 59.02
Sparse Semi-DETR [6] CVPR 24 48.34 55.26 59.19
SOOD [53] CVPR 23 48.63 55.58 59.23
SOOD++ (ours) - 50.48(+1.85) 57.44(+1.86) 61.51(+2.28)
@@
DOTA-V1.0 [70] is an early version of the DOTA-series datasets, which contains 15 categories identical to DOTA-V1.5, except for annotating container-crane. For the partially labeled data setting, we randomly sample 10%, 20%, and 30% of the images from DOTA-V1.0-train as labeled data and set the remaining images as unlabeled data. For the fully labeled data setting, we set DOTA-V1.0-train as labeled data and DOTA-V1.0-test as unlabeled. The mAP is used as the evaluation metric.
DOTA-V1.5 [71] DOTA-V1.5 contains more small instances (less than 10 pixels), which makes it more challenging compared to DOTA-V1.0. For the partially labeled data setting, we randomly sample 10%, 20%, and 30% of the images from DOTA-V1.5-train as labeled data and set the remaining images as unlabeled data. For the fully labeled data setting, we set DOTA-V1.5-train as labeled data and DOTA-V1.5-test as unlabeled data. The mAP is used as the evaluation metric.
DOTA-V2.0 [71] is an enhanced version of DOTA-V1.0, incorporating larger images, additional object categories, and comprehensive annotations of tiny instances. The resulting dataset comprises 11,268 images, 1,793,658 object instances, and 18 categories in total. For the partially labeled data setting, we randomly sample 10%, 20%, and 30% of the images from DOTA-V2.0-train as labeled data and set the remaining images as unlabeled data. For the fully labeled data setting, we set DOTA-V2.0-train as labeled data and DOTA-V2.0-test as unlabeled data. The mAP is used as the evaluation metric.
@@
HRSC2016 [72] is a widely-used dataset for arbitrarily oriented ship detection. For the partially labeled data setting, we randomly sample 10%, 20%, and 30% images from the training set as labeled data and set the remaining images as unlabeled data. For the fully labeled data setting, we set the training set as labeled data and the test set as unlabeled data. We report the AP85 and mAP50:95 performance on the validation set.
DIOR-R[43] is an oriented bounding box extension of the original DIOR dataset[73]. For the partially labeled data setting, we randomly sample 10%, 20%, and 30% of the training images as labeled data, treating the remaining images as unlabeled. For the fully labeled data setting, we use the training set as labeled data and the test set as unlabeled data. The performance is reported under the validation set.
@@
nuScenes [74] is a large-scale oriented 3D object detection dataset containing 1,000 distinct scenarios and 1.4M oriented 3D bounding boxes from 10 categories. These scenarios are recorded with a 32-beam LiDAR, six surround-view cameras, and five radars. Each scene has 20s video frames and is fully annotated with 3D bounding boxes every 0.5 seconds. For the partially labeled data setting, we randomly sample 1/12, 1/6, and 1/3 annotated frames from the training set as labeled data and set the remaining training frames as unlabeled data. For the fully labeled data setting, we set all annotated frames from the training set as labeled data and unannotated frames as unlabeled data. We use the mAP and NDS as the primary metrics. Higher values indicate better performance and the detailed information can be found in  [74].
@@ 5.2 Implementation Details
For the DOTA-V1.0, DOTA-V1.5, DOTA-V2.0, HRSC2016, and DIOR-R datasets, we take FCOS [56] as the representative anchor-free detector (the ResNet-50 [75] with FPN [76] used as the backbone), where BEVDet [77] is adopted for the nuScenes dataset. Following previous works [70, 78, 47], we crop the original images into $\times1,024$ patches ($\times800$ for DIOR-R dataset). We utilize asymmetric data augmentation for unlabeled data. We use strong augmentation (random flipping, color jittering, grayscale, and Gaussian blur) for the student model and weak augmentation (random flipping) for the teacher model. The models are trained for 180k iterations on 2 $\times$ RTX3090 GPUs (except the nuScenes dataset, using 8 $\times$ V100). We set the loss balance weight $\psi$ as 50. With the SGD optimizer, the initial learning rate of 0.0025 is divided by ten at 120k and 160k. The momentum and the weight decay are set to 0.9 and 0.0001, respectively. Each GPU takes three images as input, where the proportion between unlabeled and labeled data is set to 1:2, where the sample ratio $\mathcal{R}$ and IOU threshold $\mathcal{T}_{h}$ are set to 0.25 and 0.1, respectively. The hyperparameter noisy intensity $\beta$ and global number threshold $\mathcal{T}_{g}$ are set to 0.3 and 150. These hyperparameters are kept the same across different datasets. Following previous SSOD works [3, 1], we use the “burn-in” strategy to initialize the teacher model.
TABLE IV: Comparison on the DOTA-V1.5 (val set) with single-scale training/testing under fully labeled data. Results before the arrow are the supervised baseline (// for rotated Faster R-CNN/FCOS/Deformable DETR).
Method Publication 100%
Unbiased Teacher [1] ICLR 21 66.12 $\xrightarrow{-1.27}$ 64.85
Soft Teacher [2] ICCV 21 66.12 $\xrightarrow{+0.28}$ 66.40
Dense Teacher [3] ECCV 22 65.46 $\xrightarrow{+0.92}$ 66.38
ARSL [5] CVPR 23 65.46 $\xrightarrow{+1.22}$ 66.68
TMR-RD-v2 [7] WACV 24 66.12 $\xrightarrow{+0.81}$ 66.93
Sparse Semi-DETR [6] CVPR 24 66.20 $\xrightarrow{+0.87}$ 67.07
SOOD [53] CVPR 23 65.46 $\xrightarrow{+2.24}$ 67.70
SOOD++ (ours) - 65.46 $\xrightarrow{+3.58}$ 69.04
5.3 Main Results
We compare our method with the state-of-the-art SSOD methods [3, 2, 1, 5] on several challenging datasets. For a fair comparison, we re-implement these methods and have carefully tuned their hyper-parameters to obtain their best performance on oriented object detectors with the same augmentation setting.
5.3.1 Results on the DOTA-V2.0 dataset
We first evaluate on the large-scale DOTA-V2.0 [71] dataset from the partially labeled setting and full label setting.
Partially labeled data. We validate SOOD++ under three partially labeled settings (10%, 20%, and 30% proportion). As illustrated in Tab. I, several critical observations can be highlighted: 1) SOOD++ consistently achieves superior performance, reaching 49.14, 54.27, and 56.90 mAP at the 10%, 20%, and 30% labeling ratios, respectively, substantially exceeding our supervised baseline (rotated-FCOS[56]) by large margins of 10.36, 7.53, and 7.28 mAP. These improvements clearly indicate that SOOD++ effectively utilizes unlabeled data, significantly boosting detection performance. 2) Compared to recent advanced semi-supervised approaches, including ARSL [5], TMR-RD-v2 [7], and Sparse Semi-DETR [6], SOOD++ achieves notable superiority across all labeling settings. Particularly, under the scenario of 10% labeled data, our method surpasses these advanced competitors by substantial margins of up to 2.90 mAP, demonstrating its remarkable ability to leverage limited labeled data in conjunction with abundant unlabeled data. 3) Compared with our previous conference version, SOOD [53], SOOD++ achieves consistent improvements of 2.15, 1.92, and 2.33 mAP under the 10%, 20%, and 30% labeling ratios, respectively, reflecting substantial technical enhancements introduced in this extended version.
Fully labeled data. The fully labeled data setting presents a more challenging scenario for semi-supervised learning. In this case, the detector is already trained on abundant annotated samples, and performance tends to saturate. Consequently, leveraging additional unlabeled data to yield further improvements becomes significantly more difficult, requiring both high-quality pseudo-labels and well-designed optimization strategies. As shown in Tab. II, SOOD++ surpasses all state-of-the-art semi-supervised object detection approaches[3, 2, 7, 6], achieving a remarkable mAP of 64.31 and outperforming the best competing method by at least 2.25 mAP. Moreover, SOOD++ even exceeds the strong fully supervised baseline by 3.62 mAP, clearly validating the effectiveness of our semi-supervised learning framework. Interestingly, we observe that Unbiased Teacher [1] experiences performance degradation when unlabeled data is introduced. We conjecture that this may be attributed to its lack of unsupervised bounding box regression loss, which is particularly critical for the precise localization required in oriented object detection.
5.3.2 Results on the DOTA-V1.5 dataset
We then evaluate our method on another large-scale, challenging oriented object detection dataset (DOTA-V1.5 [71]).
Partially labeled data. Tab. III presents the results of various partially labeled data settings on the DOTA-V1.5 dataset. Specifically, the proposed SOOD++ achieves state-of-the-art performance under all proportions, achieving 50.48, 57.44, and 61.51 mAP on 10%, 20%, and 30% proportions, respectively, significantly surpassing our supervised baseline rotated-FCOS [56] by 7.70, 7.33, and 6.72 mAP. Compared with the state-of-the-art anchor-free method ARSL [5] and our conference version [53], we significantly outperform them by 2.92/1.85, 2.39/1.86, and 2.57/2.28 mAP, highly demonstrating the substantial improvements we made over the conference version. Compared with TMR-RD-v2 [7] and Sparse Semi-DETR [6], on 10% and 20% proportions, our SOOD++ achieves higher performance even though our supervised baseline (rotated-FCOS) is weaker than they used (rotated-Faster R-CNN and rotated Deformable DETR). Under 30% data proportion, SOOD++ surpasses TMR-RD-v2 [7] and Sparse Semi-DETR [6] by a large margin, i.e., 2.49 and 2.32 mAP, respectively.
Fully labeled data. As shown in Tab. IV, the results indicate that we surpass previous SSOD methods [3, 2, 7, 6] by at least 1.97 mAP. Besides, compared with our baseline, even though it has been trained well on massive labeled data, we still achieve 3.58 mAP improvement, demonstrating our method’s ability to learn from unlabeled data.
Figure 6: Comparison on the DOTA-V1.0 (val set) under 10%, 20%, and 30% partially labeled data settings, using single-scale training and testing. // indicate implementations using rotated Faster R-CNN/FCOS/Deformable DETR.
@@ 5.3 Main Results
Supervised-baseline Rotated-Faster R-CNN [32] NeurIPS 15 45.98 54.37 57.81
Rotated-FCOS [56] ICCV 19 46.20 53.88 57.30
Rotated Defor. DETR [69] ICLR 21 46.51 54.01 57.39
Semi-supervised Dense Teacher [3] ECCV 22 50.87 58.02 60.91
ARSL [5] CVPR 23 52.89 59.84 62.92
TMR-RD-v2 [7] WACV 24 52.81 59.90 63.12
Sparse Semi-DETR [6] CVPR 24 53.11 60.01 63.24
SOOD++ (ours) - 54.17 60.53 64.93
Figure 7: Comparison on the DOTA-V1.0 (val set) with single-scale training/testing under fully labeled data. Results before the arrow are the supervised baseline (// for rotated Faster R-CNN/FCOS/Deformable DETR).
Method Publication 100%
Dense Teacher [3] ECCV 22 70.22 $\xrightarrow{+0.98}$ 71.20
ARSL [5] CVPR 23 70.22 $\xrightarrow{+1.41}$ 71.63
TMR-RD-v2 [7] WACV 24 70.08 $\xrightarrow{+1.60}$ 71.68
Sparse Semi-DETR [6] CVPR 24 70.29 $\xrightarrow{+1.56}$ 71.85
SOOD++ (ours) - 70.22 $\xrightarrow{+2.18}$ 72.40
Figure 8: Comparison on the HRSC2016 (val set) under 10%, 20%, and 30% partially labeled data settings, using single-scale training and testing. // indicate implementations using rotated Faster R-CNN/FCOS/Deformable DETR. We report the AP85 and mAP50:95.
@@
AP mAP AP mAP AP mAP
Supervised-baseline Rotated-Faster R-CNN [32] NeurIPS 15 8.4 38.0 30.1 52.4 38.1 59.5
Rotated-FCOS [56] ICCV 19 7.3 37.5 30.2 52.6 37.4 58.8
Rotated Defor. DETR [69] ICLR 21 8.0 37.9 30.5 52.8 37.5 58.8
Semi-supervised Dense Teacher [3] ECCV 22 21.9 49.9 32.8 59.6 41.4 63.1
ARSL [5] CVPR 23 21.5 49.2 33.6 59.8 41.1 63.4
TMR-RD-v2 [7] WACV 24 21.4 50.1 34.4 60.1 41.9 63.5
Sparse Semi-DETR [6] CVPR 24 22.0 50.4 35.2 60.4 42.0 63.8
SOOD++ (ours) - 22.9 52.6 40.3 62.1 47.8 65.3
Figure 9: Comparison on the HRSC2016 (val set) with single-scale training/testing under fully labeled data. Results before the arrow are the supervised baseline (// for rotated Faster R-CNN/FCOS/Deformable DETR).
Method Publication 100%
Dense Teacher [3] ECCV 22 49.5 $\xrightarrow{+3.2}$ 52.7
ARSL [5] CVPR 23 49.5 $\xrightarrow{+5.0}$ 54.5
TMR-RD-v2 [7] WACV 24 49.7 $\xrightarrow{+5.3}$ 55.0
Sparse Semi-DETR [6] CVPR 24 49.6 $\xrightarrow{+5.4}$ 55.3
SOOD++ (ours) - 49.5 $\xrightarrow{+8.9}$ 58.4
5.3.3 Results on the DOTA-V1.0 dataset
We then conduct the experiments on the DOTA-V1.0 [70] dataset, which has the same images as the DOTA-V1.5 [71] dataset but provides coarse annotations (e.g., many tiny objects with fewer pixels are not annotated).
Partially labeled data. We evaluate our method under different labeled data proportions on the DOTA-V1.0 dataset, as shown in Tab. 7. It is similar to the comparison of the DOTA-V2.0 and DOTA-V1.5 datasets, where our SOOD++ achieves state-of-the-art performance under all proportions. Specifically, it surpasses the Sparse Semi-DETR [6] by 1.06, 0.52, and 1.69 mAP under 10%, 20%, and 30% labeled data settings, respectively. These impressive results demonstrate that our method handles the coarse annotation dataset well.
Fully labeled data. Under this setting, we can explore the potential of the proposed semi-supervised framework even if it has been trained with extensive labeled data. As shown in Tab. 7, although the supervised baseline has already achieved high performance, our SOOD++ still achieves 2.18 mAP improvements on it. In addition, our method performs better compared with other SOTA semi-supervised methods [6, 7].
Figure 10: Some typical examples from the DOTA-V1.5 dataset. We choose the dense pseudo-labeling (DPL)-based methods for visualization. From left to right, there are ground truth, the supervised baseline, a SOTA DPL-based method ARSL [5], and our proposed SOOD++. The red dashed and solid red circles represent false negative and false positive, respectively.
Figure 11: Comparison on the DIOR-R (val set) under 10%, 20%, and 30% partially labeled data settings, using single-scale training and testing. // indicate implementations using rotated Faster R-CNN/FCOS/Deformable DETR.
@@
Supervised-baseline Rotated-Faster R-CNN [32] NeurIPS 15 47.1 50.7 55.4
Rotated-FCOS [56] ICCV 19 46.2 49.8 54.3
Rotated Deformable DETR [69] ICLR 21 46.3 50.2 55.7
Semi-supervised Dense Teacher [3] ECCV 22 51.6 53.7 57.4
ARSL [5] CVPR 23 52.0 55.9 59.2
TMR-RD-v2 [7] WACV 24 53.2 57.3 59.5
Sparse Semi-DETR [6] CVPR 24 53.3 57.0 59.8
SOOD++(ours) - 56.2 60.1 63.1
Figure 12: Comparison on the DIOR-R (val set) with single-scale training/testing under fully labeled data. Results before the arrow are the supervised baseline (// for rotated Faster R-CNN/FCOS/Deformable DETR).
Method Publication 100%
Dense Teacher [3] ECCV 22 65.7 $\xrightarrow{+1.1}$ 66.8
ARSL [5] CVPR 23 65.7 $\xrightarrow{+2.3}$ 68.0
TMR-RD-v2 [7] WACV 24 66.1 $\xrightarrow{+2.4}$ 68.5
Sparse Semi-DETR [6] CVPR 24 65.8 $\xrightarrow{+2.5}$ 68.3
SOOD++ (ours) - 65.7 $\xrightarrow{+4.0}$ 69.7